# Code Security Vulnerability Repair Using Reinforcement Learning with Large Language Models

### 13 Jan 2024 v1

@@ Abstract
With the recent advancement of Large Language Models (LLMs), generating functionally correct code has become less complicated for a wide array of developers. While using LLMs has sped up the functional development process, it poses a heavy risk to code security. Code generation with proper security measures using LLM is a significantly more challenging task than functional code generation. Security measures may include adding a pair of lines of code with the original code, consisting of null pointer checking or prepared statements for SQL injection prevention. Currently, available code repair LLMs generate code repair by supervised fine-tuning, where the model looks at cross-entropy loss. However, the original and repaired codes are mostly similar in functionality and syntactically, except for a few (1-2) lines, which act as security measures. This imbalance between the lines needed for security measures and the functional code enforces the supervised fine-tuned model to prioritize generating functional code without adding proper security measures, which also benefits the model by resulting in minimal loss. Therefore, in this work, for security hardening and strengthening of generated code from LLMs, we propose a reinforcement learning-based method for program-specific repair with the combination of semantic and syntactic reward mechanisms that focus heavily on adding security and functional measures in the code, respectively.
@@ Introduction
Figure 1: An illustrative example showcases how the Cross-Entropy (CE) emphasizes only functionality and neglects security. Furthermore, we demonstrate that the BERT_Score and CodeBLEU scores show an improvement in comparing functionality and security measures.
The wide use of large language models further enhances the security vulnerability. Their functional correctness is heavily challenged by the potential introduction of security issues within the generated code (Sandoval et al. 2023; Pearce et al. 2022b). Despite the democratization of coding, providing increased accessibility and productivity among developers from code generated by large language models frequently falls short of established software security standards, possibly harboring vulnerabilities in approximately 40% of cases (Pearce et al. 2022a). This security challenge extends beyond individual models, with various evaluations revealing that other cutting-edge LMs (Nijkamp et al. 2023b; Fried et al. 2022), similar to Copilot, display identical security issues, as highlighted in (Li et al. 2023). A separate investigation (Khoury et al. 2023) identified that ChatGPT, in 16 out of 21 security-relevant cases, generates code falling below minimal security standards. The more alarming concern is the training data of generative AI models, deeply rooted in the same insecure public repositories (Chen et al. 2021). As a result, unless pre-trained explicitly in security, generative AI models give rise to a vicious cycle of generating functional code by disseminating vulnerability.
@@
Furthermore, functional code generation is a task where the model has to generate entire code based on specific instructions. However, Code1 and Code2 in Figure 1 demonstrate that security repair merely adds a few lines to the code. Here, Code 1 is the ground truth repaired function, and Code 2 is the model-generated outcome that attempts to repair a vulnerable function. Original input to the model is the exclusion of the line dest[99] = NULL. From the ground truth and the generated repair by the model, we observe that the cross-entropy (CE) loss between the original and model outcomes is very low. Thus, the model is tempted to ignore the security measures and generate code almost similar to the input since it yields minimum loss. Moreover, BERTScore on Code1 and Code2 shows higher loss values. Similarly, if we only compare the security measure with a syntactic similarity-based measurement CodeBLEU, we still see that the loss value remains much higher. These higher values indicate that cross-entropy loss is insufficient to make the generative model security aware.

### 30 Jan 2024 v2

@@ Abstract
With the recent advancement of Large Language Models (LLMs), generating functionally correct code has become less complicated for a wide array of developers. While using LLMs has sped up the functional development process, it poses a heavy risk to code security. Code generation with proper security measures using LLM is a significantly more challenging task than functional code generation. Security measures may include adding a pair of lines of code with the original code, consisting of null pointer checking or prepared statements for SQL injection prevention. Currently, available code repair LLMs generate code repair by supervised fine-tuning, where the model looks at cross-entropy loss. However, the original and repaired codes are mostly similar in functionality and syntactically, except for a few (1-2) lines, which act as security measures. This imbalance between the lines needed for security measures and the functional code enforces the supervised fine-tuned model to prioritize generating functional code without adding proper security measures, which also benefits the model by resulting in minimal loss. Therefore, in this work, for security hardening and strengthening of generated code from LLMs, we propose a reinforcement learning-based method for program-specific repair with an attempt to combine semantic and syntactic reward mechanisms that focus heavily on adding security and functional measures in the code, respectively.
@@ Introduction
Figure 1: An illustrative example showcases how the Cross-Entropy (CE) emphasizes only functionality and neglects security.
The wide use of large language models further enhances the security vulnerability. Their functional correctness is heavily challenged by the potential introduction of security issues within the generated code (Sandoval et al. 2023; Pearce et al. 2022b). Despite the democratization of coding, providing increased accessibility and productivity among developers from code generated by large language models frequently falls short of established software security standards, possibly harboring vulnerabilities in approximately 40% of cases (Pearce et al. 2022a). This security challenge extends beyond individual models, with various evaluations revealing that other cutting-edge LMs (Nijkamp et al. 2023b; Fried et al. 2022), similar to Copilot, display identical security issues, as highlighted in (Li et al. 2023). A separate investigation carried out by (Khoury et al. 2023) identified that ChatGPT, in 16 out of 21 security-relevant cases, generates code falling below minimal security standards. The more alarming concern is the training data of generative AI models, deeply rooted in the same insecure public repositories (Chen et al. 2021). As a result, unless pre-trained explicitly in security, generative AI models give rise to a vicious cycle of generating functional code by disseminating vulnerability.
@@
Furthermore, functional code generation is a task where the model has to generate entire code, which is based on specific instructions. The functional task of Code 1 and Code 2 is to copy data from the source to the destination. However, Code1 and Code 2 in Figure 1 demonstrate that security repair merely adds a few lines to the code. Here, Code 1 is the ground truth repaired function, and Code 2 is the model-generated outcome that attempts to repair a vulnerable function. The original vulnerable input to the model is the exclusion of the line dest[99] = NULL. From the ground truth. We observe that the cross-entropy (CE) loss between the original and model outcomes is very low. Thus, the model is tempted to ignore the security measures and generate code almost similar to the input since it yields minimum loss. This issue effectively demonstrates that the use of CE loss only provides us with token-based similarity when we compare the generated code with the ground truth. However, this also demonstrates that merely checking code similarity without checking the semantics and syntactic does not do a proper evaluation of the code.
@@ Discussion:
In this section, we will provide three case studies which will compare the outcome of our RL model vs. the SFT model.
@@ 6    Memcached *root = (Memcached *) instance->root;
Threats to Validity
While we see an improvement in the performance when using RL for vulnerability repair, the methods we use to evaluate the code are not designed for code comparison or understanding the semantics of code security. BLEU and Rouge-L metrics only does a token based comparison, and does not adhere to the security components of the repaired code. Furthermore, our case studies shows that our model was able to generate code with security repairs, these metrics does not reflect the performance. Therefore, in order to do code comparison, and security evaluation of source code, we need proper measurements which will comply with not only the functionality of code but also with the added security measurements of the code.